# Breaking Imitation Bottlenecks: Reinforced Diffusion Powers Diverse Trajectory Generation

### v2

@@ 3.2 Diffusion Models for Driving Policy Diversity
Figure 2: Imitation learning-based multi-mode trajectories paradigm. Most IL-based multi-mode E2E-AD methods rely on L1 loss for training and L2 distance for evaluation, which emphasizes matching a single GT trajectory rather than modeling diversity. This misalignment limits the generation of truly diverse behaviors. Even with diffusion-based frameworks [4], such imitation-driven objectives constrain their capacity to capture multi-mode driving patterns.
Under conventional imitation learning (IL) settings, where training is supervised by a single ground-truth (GT) trajectory, their ability to generate diverse behaviors is fundamentally constrained. We present a theoretical analysis to explain this limitation.
Let a trajectory be defined as a sequence of waypoints $\tau={(x_{t},y_{t})}{t=1}^{T_{f}}$, and a multi-mode output as ${\tau^{(m)}}{m=1}^{M}$. In IL-based trajectory prediction, the training objective typically maximizes the log-likelihood of the observed trajectory:
$\max_{\theta}\mathbb{E}_{\tau\sim p_{\text{data}}}\left[\log p_{\theta}(\tau)\right]$ (1)
@@
$p_{\text{data}}(\tau)=\delta(\tau-\tau^{*})$ (2)
@@
$p_{\theta}^{\text{opt}}(\tau)=\delta(\tau-\tau^{*})$ (4)
Diffusion models are trained by minimizing a denoising score-matching loss, often expressed as:
$\min_{\theta}\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}\left[\left\|\epsilon-\epsilon_{\theta}(\tau^{*}+\sqrt{\alpha_{t}}\epsilon,t)\right\|^{2}\right]$ (5)
This formulation teaches the model to reconstruct $\tau^{*}$ from noisy perturbations, effectively collapsing all modes toward the single GT trajectory. No matter how strong the underlying generative process is, the diversity of output trajectories ${\tau^{(m)}}_{m=1}^{M}$collapses around $\tau^{*}$, failing to reflect genuine multi-modeity.
@@
$\Sigma=\mathbb{E}_{\tau\sim p_{\theta}}\left[(\tau-\mu)(\tau^{*}-\mu)^{T}\right],\quad\mu=\mathbb{E}[\tau]$ (6)
With all $\tau$ approximating $\tau^{*}$, we have $\mu\approx\tau^{*}$ and $\Sigma\approx 0$, indicating near-zero diversity across predicted trajectories.
In summary, diffusion-based E2E-AD methods trained under single-GT supervision inherit the same mode collapse issue as standard IL, despite their generative capacity. To address this, we propose a hybrid IL-RL training framework, DIVER, that introduces diversity-aware rewards to encourage multi-mode, safe trajectories generation.
Figure 3: The overall architecture of DIVER. As a multi-mode trajectories E2E-AD framework, DIVER first encodes multi-view images into feature maps to extract scene representations through a perception module. It then predicts the motion of surrounding agents and performs planning via a conditional diffusion model guided by reinforcement learning to generate diverse multi-intention trajectories. Our approach effectively addresses the inherent mode collapse in imitation learning, enabling the generation of safe and diverse behaviors for complex driving scenarios.
@@ 4.1 Overall Architecture
As shown in Figure 3, we propose DIVER, a novel E2E-AD framework that integrates diffusion models and reinforcement learning to enable diverse, and safe multi-mode trajectories. DIVER consists of two main components: perception module and motion planner, which takes raw perception features as input and outputs a distribution over future trajectories. At the training paradigm level, DIVER adopts a hybrid learning scheme that combines IL with RL. The core innovations of DIVER lie in two aspects. First, the Policy-Aware Diffusion Generator (PADG) generates diverse trajectory candidates by leveraging a diffusion model conditioned not only on a single expert trajectory but also on multiple Reference GTs, capturing a range of plausible human-like driving behaviors such as lane changes, yielding, and overtaking. Second, we leverage reinforcement learning to guide the diffusion-based trajectory generation process, addressing the inherent limitations of IL-based supervision and aligning the output with safety and diversity objectives. Overall, DIVER combines the generative capability of diffusion models with the optimization strength of reinforcement learning, forming a flexible and scalable framework for E2E-AD. It enables the system to explore a broader solution space in complex scenarios while ensuring the physical feasibility and behavioral safety of the generated trajectories.
4.2 Policy-Aware Diffusion Generator.
@@
4.2.1 Diverse Multi-Mode Trajectories Forward Diffusion.
To generate diverse and informative future trajectory distributions, we apply a forward diffusion process to multiple trajectory types, including multi-mode predicted trajectories, anchor proposals, and the GT trajectory. This step aims to perturb the trajectories with Gaussian noise in a controlled manner, enabling the conditional diffusion model to learn a robust denoising process that captures the underlying multi-modality of future motion.
Specifically, let $\tau_{0}^{(m)}\in\mathbb{R}^{T\times 2}$ denote a predicted future trajectory of mode $m$, $\tau_{0}^{\text{a}(m)}$ denote its associated anchor proposal, and $\tau_{0}^{\text{gt}}$ the GT future trajectory. At each diffusion step $t=1,\dots,T$, the trajectories are corrupted with Gaussian noise following the standard forward diffusion process:
@@
$q\left(\tau^{\text{a}(m)}_{t}\mid\tau^{\text{a}(m)}_{0}\right)$ $=\mathcal{N}\left(\tau^{\text{a}(m)}_{t};\sqrt{\alpha_{t}}\tau^{\text{a}(m)}_{0},(1-\alpha_{t})\mathbf{I}\right),$ (8)
$q\left(\tau^{\text{gt}}_{t}\mid\tau^{\text{gt}}_{0}\right)$ $=\mathcal{N}\left(\tau^{\text{gt}}_{t};\sqrt{\alpha_{t}}\tau^{\text{gt}}_{0},(1-\alpha_{t})\mathbf{I}\right)$ (9)
where $\alpha_{t}\in(0,1)$ is the noise schedule at timestep $t$. This formulation ensures that the model can learn to recover high-quality trajectory samples from various noise levels during training, while preserving the structured correlations between predicted and anchor trajectories.
By jointly applying noise to both predicted and anchor trajectories, we encourage the model to explore the latent trajectory space more broadly, leading to greater trajectory diversity during generation. Simultaneously, injecting noise into the GT trajectory supports supervised learning of the denoising objective, grounding the generative process in realistic motion patterns.
Figure 4: The illustration of Policy-Aware Diffusion Generator. By incorporating the predicted trajectory, GT trajectory, and anchor trajectory as inputs, PADG reconstructs diverse multi-mode trajectories from noise through a conditional denoising process, guided by map and agent context.
4.2.2 Conditional Diffusion Decoder.
@@
$\mathcal{E}^{(m)}{\tau}$ $=\text{PE}(\tilde{\tau}^{(m)}),\mathcal{F}^{(m)}{\tau}=\text{Enc}(\mathcal{E}^{(m)}{\tau}),$ (10)
$\mathcal{E}^{(m)}{\tau^{\text{a}}}$ $=\text{PE}(\tilde{\tau}^{\text{a}(m)}),\mathcal{F}^{(m)}{\tau^{\text{a}}}=\text{Enc}(\mathcal{E}^{(m)}{\tau^{\text{a}}}),$ (11)
@@
Given the GT trajectory $\tau^{\text{GT}}$ and its diffusion noise $\epsilon\sim\mathcal{N}(0,I)$, we obtain noisy trajectories using the forward process of the diffusion scheduler:
$\tilde{\tau}_{t}^{(m)}=q_{t}(\tau^{\text{GT}},\epsilon)=\sqrt{\bar{\alpha}_{t}}\tau^{\text{GT}}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ (12)
where $t$ denotes the diffusion step and $\bar{\alpha}_{t}$ the cumulative noise factor. We first embed both the noisy predicted trajectory $\tilde{\tau}_{t}^{(m)}$ and the corresponding reference trajectory $\tau^{\text{ref}(m)}$ to extract their spatial-temporal semantic features. The embedding process is defined as:
$\mathcal{F}_{\tau}^{(m)}$ $=\text{PE}\left(\phi\left(\tilde{\tau}t^{(m)}\right)\right),$ (13)
$\mathcal{F}{\text{ref}}^{(m)}$ $=\text{PE}\left(\phi\left(\tau^{\text{ref}(m)}\right)\right),$ (14)
where $\phi(\cdot)$ denotes the sine-cosine positional embedding applied to each trajectory waypoint, i.e., $\phi(\cdot)=\text{SineEmbed}(\cdot)$, $\text{PE}(\cdot)$ denotes the trajectory position encoder, typically implemented as a lightweight MLP or transformer block, $\mathcal{F}{\tau}^{(m)}$ and $\mathcal{F}{\text{ref}}^{(m)}$ are the learned representations of the noisy predicted trajectory and the reference trajectory for mode $m$, respectively.
@@
In the first stage, we apply a trajectory-aware image feature pooling module, referred to as TrajPooler (Algorithm 1). Given a noisy trajectory $\hat{\tau}^{(m)}$, we first recover its absolute coordinates via cumulative sum. We then sparsely sample 3D points along the trajectory and project them onto the image plane using the known camera intrinsics $\mathbf{K}$ and extrinsics $\mathbf{T}$. These projected 2D locations are used to extract multi-scale image features from the BEV encoder’s output. To adaptively aggregate these image features, we learn attention weights $\alpha^{(m)}$ conditioned on the initial trajectory feature $F{\text{traj}}^{(m)}$. The fused trajectory feature $F{\text{traj}}^{(m)}$ is obtained via a weighted sum followed by an MLP and residual fusion with the original instance embedding.
In the second stage, we inject the pooled trajectory token into a transformer-based conditional decoder, as shown in Algorithm 2. Specifically, we follow a hierarchical attention stack where the trajectory token $F_{\text{traj}}^{(m)}$ is first used to query agent-level memory via multi-head cross-attention (MHCA), producing agent-conditioned features $F_{\text{agent}}^{(m)}$. These are further refined by attending to map-level memory features, yielding $F_{\text{map}}^{(m)}$. Finally, navigation-level anchors (e.g., waypoints from reference plans or diffusion priors) are fused via another MHCA block. The final output $F_{\text{out}}^{(m)}$ is obtained by feeding the resulting feature through a residual FFN block.
@@
Trajectory token $F_{\text{traj}}^{(m)}\in\mathbb{R}^{d}$,
Agent memory tokens $\mathcal{F}_{\text{agent}}\in\mathbb{R}^{N_{a}\times d}$ with $\mathbf{P}_{\text{agent}}$,
Map memory tokens $\mathcal{F}_{\text{map}}\in\mathbb{R}^{N_{m}\times d}$ with $\mathbf{P}_{\text{map}}$,
Navigation memory $\mathcal{F}_{\text{nav}}^{(m)}\in\mathbb{R}^{T\times d}$,
@@
2Agent-Level Cross Attention: $F_{\text{agent}}^{(m)}=\text{MHCA}(Q^{(m)}+F_{\text{traj}}^{(m)},\ \mathcal{F}_{\text{agent}}+\mathbf{P}_{\text{agent}},\ \mathcal{F}_{\text{agent}})$
3Map-Level Cross Attention: $F_{\text{map}}^{(m)}=\text{MHCA}(Q^{(m)}+F_{\text{agent}}^{(m)},\ \mathcal{F}_{\text{map}}+\mathbf{P}_{\text{map}},\ \mathcal{F}_{\text{map}})$
4Anchor-Level Cross Attention: $F_{\text{nav}}^{(m)}=\text{MHCA}(F_{\text{map}}^{(m)},\ \mathcal{F}_{\text{nav}}^{(m)},\ \mathcal{F}_{\text{nav}}^{(m)})$
Final LayerNorm + FFN: $F_{\text{out}}^{(m)}=\text{FFN}(\text{LayerNorm}(F_{\text{nav}}^{(m)})+F_{\text{nav}}^{(m)})$
@@
At inference time, the model samples multiple noise vectors $\epsilon^{(m)}\sim\mathcal{N}(0,I)$ and generates corresponding multi-mode trajectories $\hat{\tau}^{(m)}$ via the denoising diffusion process. The reference trajectories $\tau^{\text{ref}(m)}$ can be drawn from anchor plans, plan queries, or diffusion priors, ensuring that the output space spans diverse yet goal-directed behaviors.
@@ 4.3.1 Overview of Group Relative Policy Optimization
$\mathcal{J}_{\mathrm{GRPO}}(\theta)=\mathbb{E}_{q,\left\{o_{i}\right\}\sim\pi_{\theta_{old}}}\left[\frac{1}{G}\sum_{i=1}^{G}\mathcal{L}_{i}-\beta\mathbb{D}_{KL}\left(\pi_{\theta}\|\pi_{\text{ref }}\right)\right],$ (16)
@@
where, $w_{i}=\frac{\pi_{\theta}\left(o_{i}\mid q\right)}{\pi_{\theta_{old}}\left(o_{i}\mid q\right)}$, $\epsilon$ and $\beta$ are hyper-parameters, and the advantage $A_{i}$ is computed using the normalized reward within the group.

### v3

@@ 3.2 Diffusion Models for Driving Policy Diversity
Figure 2: Imitation learning-based multi-mode trajectory paradigm. Most IL-based multi-mode E2E-AD methods rely on L1 loss for training and L2 distance for evaluation, which emphasizes matching a single GT trajectory rather than modeling diversity. This misalignment limits the generation of truly diverse behaviors. Even with diffusion-based frameworks [8], such imitation-driven objectives constrain their capacity to capture multi-mode driving patterns.
Under conventional imitation learning (IL) settings, where training is supervised by a single ground-truth (GT) trajectory, their ability to generate diverse behaviors is fundamentally constrained. We present a theoretical analysis to explain this limitation. Let a trajectory be defined as a sequence of waypoints $\tau={(x_{t},y_{t})}_{t=1}^{T_{f}}$, and a multi-mode output as $\{\tau^{(m)}\}_{m=1}^{M}$. In IL-based trajectory prediction, the training objective typically maximizes the log-likelihood of the observed trajectory:
$\max_{\theta}\mathbb{E}_{\tau\sim p_{\text{data}}}\left[\log p_{\theta}(\tau)\right].$ (1)
@@
$p_{\text{data}}(\tau)=\delta(\tau-\tau^{*}).$ (2)
@@
$p_{\theta}^{\text{opt}}(\tau)=\delta(\tau-\tau^{*}).$ (4)
In contrast, diffusion models are trained by minimizing a denoising score-matching loss, often expressed as:
$\min_{\theta}\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}\left[\left\|\epsilon-\epsilon_{\theta}(\tau^{*}+\sqrt{\alpha_{t}}\epsilon,t)\right\|^{2}\right].$ (5)
This formulation teaches the model to reconstruct $\tau^{*}$ from noisy perturbations, effectively collapsing all modes toward the single GT trajectory. No matter how strong the underlying generative process is, the diversity of output trajectories $\{\tau^{(m)}\}_{m=1}^{M}$collapses around $\tau^{*}$, failing to reflect genuine multi-modality.
@@
$\Sigma=\mathbb{E}_{\tau\sim p_{\theta}}\left[(\tau-\mu)(\tau^{*}-\mu)^{T}\right],\quad\mu=\mathbb{E}[\tau].$ (6)
With all $\tau$ approximating $\tau^{*}$, we have $\mu\approx\tau^{*}$ and $\Sigma\approx 0$, indicating near-zero diversity across predicted trajectories. In summary, diffusion-based E2E-AD methods trained under single-GT supervision inherit the same mode collapse issue as standard IL, despite their generative capacity. To address this, we propose a hybrid IL-RL training framework, DIVER, that introduces diversity-aware rewards to encourage multi-mode, safe trajectories generation.
Figure 3: The overall architecture of DIVER. As a multi-mode trajectory E2E-AD framework, DIVER first encodes multi-view images into feature maps to extract scene representations through a perception module. It then predicts the motion of surrounding agents and performs planning via a conditional diffusion model guided by reinforcement learning to generate diverse multi-mode trajectories. Our approach effectively addresses the inherent mode collapse in imitation learning, enabling the generation of safe and diverse behaviors for complex driving scenarios.
@@ 4.1 Overall Architecture
Existing E2E-AD methods [1, 2, 6, 8, 3] predominantly rely on imitation learning (IL) from single expert demonstration, which often leads to conservative and homogeneous driving behaviors, thereby limiting their generalization to complex real-world scenarios. Although some diffusion-based methods, such as DiffusionDrive [8], introduce diffusion models to mitigate mode collapse, their training remains constrained by imitation learning and lacks explicit optimization for task-level objectives such as safety or lane compliance. Therefore, as shown in Figure 3, we propose DIVER, a novel E2E-AD framework that integrates diffusion models and reinforcement learning to enable diverse, and safe multi-mode trajectories. DIVER consists of two main components: perception module and motion planner, which takes raw perception features as input and outputs a distribution over future trajectories. At the training paradigm level, DIVER adopts a hybrid learning scheme that combines IL with RL. The core innovations of DIVER lie in two aspects. First, the Policy-Aware Diffusion Generator (PADG) generates diverse trajectory candidates by leveraging a diffusion model conditioned not only on a single expert trajectory but also on multiple Reference GTs, capturing a range of plausible human-like driving behaviors such as lane changes, yielding, and overtaking. Second, we leverage reinforcement learning to guide the diffusion-based trajectory generation process, addressing the inherent limitations of IL-based supervision and aligning the output with safety and diversity objectives. Overall, DIVER combines the generative capability of diffusion models with the optimization strength of reinforcement learning, forming a flexible and scalable framework for E2E-AD. It enables the system to explore a broader solution space in complex scenarios while ensuring the physical feasibility and behavioral safety of the generated trajectories.
4.2 Policy-Aware Diffusion Generator
@@
4.2.1 Diverse Multi-Mode Trajectory Forward Diffusion
To generate diverse and informative future trajectory distributions, we apply a forward diffusion process to multiple trajectory types, including multi-mode predicted trajectories and the GT trajectory. This step aims to perturb the trajectories with Gaussian noise in a controlled manner, enabling the conditional diffusion model to learn a robust denoising process that captures the underlying multi-modeity of future motion.
Specifically, let $\tau_{0}^{(m)}\in\mathbb{R}^{T\times 2}$ denote a predicted future trajectory of mode $m$, and $\tau_{0}^{\text{gt}}$ the GT future trajectory. At each diffusion step $t=1,\dots,T$, the trajectories are corrupted with Gaussian noise following the standard forward diffusion process:
@@ 4.2.1 Diverse Multi-Mode Trajectory Forward Diffusion
$q\left(\tau^{\text{gt}}_{t}\mid\tau^{\text{gt}}_{0}\right)$ $=\mathcal{N}\left(\tau^{\text{gt}}_{t};\sqrt{\alpha_{t}}\tau^{\text{gt}}_{0},(1-\alpha_{t})\mathbf{I}\right),$ (8)
where $\alpha_{t}\in(0,1)$ is the noise schedule at timestep $t$. This formulation ensures that the model can learn to recover high-quality trajectory samples from various noise levels during training, while preserving the structured correlations of the predicted trajectories.
By jointly applying noise to predicted trajectories, we encourage the model to explore the latent trajectory space more broadly, leading to greater trajectory diversity during generation. Simultaneously, injecting noise into the GT trajectory supports supervised learning of the denoising objective, grounding the generative process in realistic motion patterns.
Figure 4: The illustration of Policy-Aware Diffusion Generator. By incorporating the predicted trajectory, and GT trajectory as inputs, PADG reconstructs diverse multi-mode trajectories from noise through a conditional denoising process, guided by map and agent context.
4.2.2 Conditional Diffusion Decoder
@@
$\mathcal{E}^{(m)}_{\tau}$ $=\text{PE}(\tilde{\tau}^{(m)}),\mathcal{F}^{(m)}_{\tau}=\text{Enc}(\mathcal{E}^{(m)}_{\tau})$, (9)
$\mathcal{E}^{(m)}_{\tau^{\text{a}}}$ $=\text{PE}(\tilde{\tau}^{\text{a}(m)}),\mathcal{F}^{(m)}_{\tau^{\text{a}}}=\text{Enc}(\mathcal{E}^{(m)}_{\tau^{\text{a}}})$, (10)
@@
Given the noisy GT trajectory $\tilde{\tau}^{\text{gt}}_{t}$ obtained from the forward diffusion process, we use it as the reference trajectory $\tau^{\text{ref}(m)}$ for feature embedding:
$\mathcal{F}_{\text{ref}}^{(m)}$ $=\text{PE}\left(\phi\left(\tau^{\text{ref}(m)}\right)\right),$ (11)
where $\phi(\cdot)$ denotes sine-cosine positional embeddings applied to each trajectory waypoint, and $\text{PE}(\cdot)$ is a trajectory position encoder. Here, $\mathcal{F}_{\text{ref}}^{(m)}$ represents the learned feature of the noisy GT trajectory, which is then used for multi-mode interaction with high-level scene features in the conditional diffusion decoder.
@@
In the first stage, we adopt a lightweight trajectory-aware feature aggregation module, termed TrajPooler (Algorithm 1), which focuses on learning perception features around the predicted trajectory rather than processing the entire image. given a noisy trajectory $\hat{\tau}^{(m)}$, we first recover its absolute coordinates via cumulative summation. We then sparsely sample 3D points along the trajectory and project them onto the image plane using the known camera intrinsics $\mathbf{K}$ and extrinsics $\mathbf{T}$. Multi-scale image features are extracted at these projected locations from the BEV encoder output. To adaptively aggregate these features, attention weights $\alpha^{(m)}$ are learned conditioned on the initial trajectory feature $F_{\text{traj}}^{(m)}$. The final trajectory-aware representation $F_{\text{traj}}^{(m)}$ is obtained through a weighted summation of the sampled features, followed by an MLP and residual fusion with the original trajectory embedding.
In the second stage, we inject the pooled trajectory token into a transformer-based conditional decoder, as shown in Algorithm 2. Specifically, we follow a hierarchical attention stack where the trajectory token $F_{\text{traj}}^{(m)}$ is first used to query agent-level memory via multi-head cross-attention (MHCA), producing agent-conditioned features $F_{\text{agent}}^{(m)}$. These are further refined by attending to map-level memory features, yielding $F_{\text{map}}^{(m)}$. Finally, navigation-level anchors (e.g., waypoints from reference plans or diffusion priors) are fused via another MHCA block. The final output $F_{\text{out}}^{(m)}$ is obtained by feeding the resulting feature through a residual FFN block. It should be emphasized that the anchor trajectories in DIVER are conceptually different from the reference candidates in VADv2 or Hydra-MDP, which are used for final trajectory selection. In contrast, DIVER employs only six anchors as generative guidance to encourage diverse motion patterns, rather than selecting the final output from them as VADv2 or Hydra-MDP does.
@@
Trajectory token $F_{\text{traj}}^{(m)}$
Agent memory tokens $\mathcal{F}_{\text{agent}}$
Map memory tokens $\mathcal{F}_{\text{map}}$
Navigation memory $\mathcal{F}^{(m)}{\tau^{\text{a}}}$
@@
2Agent-Level Cross Attention:
[…]
4Map-Level Cross Attention: $F_{\text{traj}}^{(m)}=\text{MHCA}(Q^{(m)}+F_{\text{traj}}^{(m)},\ \mathcal{F}_{\text{map}},\ \mathcal{F}_{\text{map}})$
5Anchor-Level Cross Attention: $F_{\text{traj}}^{(m)}=\text{MHCA}(F_{\text{traj}}^{(m)},\mathcal{F}^{(m)}{\tau^{\text{a}}},\mathcal{F}^{(m)}{\tau^{\text{a}}}$
Final LayerNorm + FFN: $F_{\text{out}}^{(m)}=\text{FFN}(\text{LayerNorm}(\mathcal{F}^{(m)}{\tau^{\text{a}}})+\mathcal{F}^{(m)}{\tau^{\text{a}}})$
@@
At inference time, the model samples multiple noise vectors $\epsilon^{(m)}\sim\mathcal{N}(0,I)$ and generates corresponding multi-mode trajectories $\hat{\tau}^{(m)}$ via the denoising diffusion process. The reference trajectories $\tau^{\text{ref}(m)}$ can be drawn from anchor plans, plan queries, or diffusion priors, ensuring that the output space spans diverse yet goal-directed behaviors. Finally, we follow the strategy of SparseDrive [6] and DiffusionDrive [8], employing a post-processing procedure in which the remaining trajectories are scored based on comfort and scene compliance, and the trajectory with the highest score is selected as the final output.
@@ 4.3.1 Overview of Group Relative Policy Optimization
$\mathcal{J}_{\mathrm{GRPO}}(\theta)=\mathbb{E}_{q,\left\{o_{i}\right\}\sim\pi_{\theta_{\textrm{old}}}}\left[\frac{1}{G}\sum_{i=1}^{G}\mathcal{L}_{i}-\beta\mathbb{D}_{KL}\left(\pi_{\theta}\|\pi_{\text{ref }}\right)\right],$ (13)
@@
where, $w_{i}=\frac{\pi_{\theta}\left(o_{i}\mid q\right)}{\pi_{\theta_{\mathrm{old}}}\left(o_{i}\mid q\right)}$, $\epsilon$ and $\beta$ are hyper-parameters, the old policy $\pi_{\theta_{\textrm{old}}}$ and the advantage $A_{i}$ is computed using the normalized reward within the group.